ON WAITING TIME PROBLEMS ASSOCIATED WITH RUNS IN MARKOV DEPENDENT TRIALS

DEMETRIOS L. ANTZOULAKOS

Department of Economics, The School of Social Sciences, University of Crete, Perivolia - Rethymno 74100, Greece

(Received July 11, 1997; revised October 13, 1997)

Abstract. A general technique is developed to study the waiting time distribution for the r-th occurrence of a success run of length k in a sequence of Markov dependent trials. Sooner and later waiting time problems are also discussed.

Key words and phrases: Probability mass function, probability generating function, waiting time, run, sooner and later problems, distributions of order k.